A bicycle moves from rest at 8.00 m/s in a time of 2.00 seconds. What is...

  1. A bicycle moves from rest at 8.00 m/s in a time of 2.00 seconds. What is its acceleration?
Answer #1

Initial speed of by cycle (v0) = 0

Time (t) = 2 sec

Final speed (v) = 8 m/s

From equation of motion -

v = v0 + at

8 = 0 + 2a

  Acceleration (a) = 4 m/s2
Answer #2

To find the acceleration of the bicycle, we can use the equation for acceleration:

Acceleration()=Change in velocity(Δ)Time()

The initial velocity () of the bicycle is 0 m/s (as it starts from rest), and the final velocity () is 8.00 m/s. The time taken () is 2.00 seconds.

Δ==8.00m/s0m/s=8.00m/s=8.00m/s2.00s=4.00m/s2

So, the acceleration of the bicycle is 4.00m/s2.


answered by: mervetokaz